OpenStudy (anonymous):

I need help. I`m trying to finish my FLVS and the equation is 8/x + 2/9 = 1/3. I know the x = 72 but can someone show me how to solve it?

OpenStudy (anonymous):

\[\frac{8}{x} + \frac{2}{9} = \frac{1}{3}\]\[\frac{8}{x} \times \frac{9}{9}+ \frac{2}{9} \times \frac{x}{x}= \frac{1}{3} \times \frac{3x}{3x}\]\[\frac{72 + 2x}{9x} = \frac{3x}{9x}\]Drop the denominators.\[72 + 2x = 3x\]\[72 = x\]

OpenStudy (anonymous):

Btw, what math are you doing on flvs?

OpenStudy (anonymous):

Algebra 2 honors!

OpenStudy (anonymous):

Oh. I literally JUST took the final exam yesterday! lol

OpenStudy (anonymous):

Haha I got another question. Can you help?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

It`s just like the one i asked but instead of us given the equations it just says x=64. So what are the equations?

OpenStudy (anonymous):

I don't see how that works. Is it an essay problem?

OpenStudy (anonymous):

Well it`s for the collaboration activity with candy.

OpenStudy (anonymous):

Oh ok. THat one. You have to basically set up a rational expression using the candy. I did that project alone...let me recheck the details.

OpenStudy (anonymous):

Okay thanks (:

OpenStudy (anonymous):

Are you trying to finish tonight? The deadline is tomorrow.

OpenStudy (anonymous):

Yes I`m trying to finish tonight.

OpenStudy (anonymous):

ALright. Yeah. all work has to be in by midnight tonight.

OpenStudy (anonymous):

Yeah I know. I just have

OpenStudy (anonymous):

*i just have to do the collaboration then two other things and I`m done.

OpenStudy (anonymous):

Alright. Did they give you the information on how much candy already?

OpenStudy (anonymous):

Or did you make it up?

OpenStudy (anonymous):

Made it up. Lol. If 64 is too hard you can change it. :p

OpenStudy (anonymous):

Alright. Did you set up rational expressions or rather numbers you got from the candy?

OpenStudy (anonymous):

Like \[\frac{8}{64} + \frac{32}{64} = \frac{40}{64}\]Then you simplified everything except the \[\frac{8}{64}\]and put x in for the 64 in what I just put above. \[\frac{8}{x} + \frac{1}{2} = \frac{5}{8}\]

OpenStudy (anonymous):

Something like that?

OpenStudy (anonymous):

Yeah so the x would equal 64 right?

OpenStudy (anonymous):

Yeah. You're supposed to give it to another student and solve for x.

OpenStudy (anonymous):

I know my friend is here with me and thats the equation I have to solve. So how would I solve it?

OpenStudy (anonymous):

\[\frac{8}{x} + \frac{1}{2} = \frac{5}{8}\]\[\frac{8}{x} \times \frac{8}{8} + \frac{1}{2} \times \frac{4x}{4x}= \frac{5}{8} \times \frac{x}{x}\]\[\frac{64}{8x} + \frac{4x}{8x} = \frac{5x}{8x}\]You can drop the denominators now that they have a common denominator. \[64 + 4x = 5x\]

OpenStudy (anonymous):

Can you're friend solve it now?

OpenStudy (anonymous):

*your

OpenStudy (anonymous):

No. lol.

OpenStudy (anonymous):

Do you understand what I was doing when I multiplied stuff?

OpenStudy (anonymous):

No I don`t. I feel stupid. Lol.

OpenStudy (anonymous):

Ok. What I'm doing there is finding a common denominator. The common denominator is 8x. What would I have to multiply x by to get 8x?

OpenStudy (anonymous):

8.

OpenStudy (anonymous):

ALright. However, if you multiplied the denominator, you have to do the same to the numerator. Do you see why I multiplied by \[\frac{8}{8}\]? I'm multiplying in both the numerator and the denominator by the same amount.

OpenStudy (anonymous):

I got it now. Thank you (: Did you do the Activity for Module 6?

OpenStudy (anonymous):

Yup. I did the rectangular prism because it has less decimals and less stuff with π in it.

OpenStudy (anonymous):

Just read the directions and it'll tell you what to do. I can't really help you much until you have your polynomial set up.

OpenStudy (anonymous):

Well I`ll try to set it up really quick if you can still help me after?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Thank you so much (: Did you take Honors or general?

OpenStudy (anonymous):

Honors

OpenStudy (anonymous):

I did Module 10 too.

OpenStudy (anonymous):

Yeah me too. Was the final hard?

OpenStudy (anonymous):

Um...there's a practice final exam to see how hard it is. It's about 53 questions with like 7 essay questions.

OpenStudy (anonymous):

Literally, anything that's not on the practice is not on the real final.

OpenStudy (anonymous):

ohh okay. if i have any questions I`m going to ask you! Lol.

OpenStudy (anonymous):

lol. Just hope that I'm online when that happens. :)

OpenStudy (anonymous):

Well I`m about to take it as soon as I finish this activity.

OpenStudy (anonymous):

Don't you have another activity?

OpenStudy (anonymous):

The activity with the box from module 6.

OpenStudy (anonymous):

I`m kind of confused on this one too. The length is 8 inches The width is 2 inches And the height is 10 inches

OpenStudy (anonymous):

Ok. THose are the measurements. What did it say to do next?

OpenStudy (anonymous):

To find the volume. So i did that and it =160

OpenStudy (anonymous):

Ok. After that?

OpenStudy (anonymous):

Now suppose you knew the volume of this object and the relation of the length to the width and height, but did not know the length. Rewriting the equation with one variable would result in a polynomial equation that you could solve to find the length. Rewrite the formula using the variable x for the length. Substitute the value of the volume found in step 2 for V and express the width and height of the object in terms of x plus or minus a constant. For example, if the height measurement is 4 inches longer than the length, then the expression for the height will be (x + 4). Simplify the equation and write it in standard form. Find the solutions to this equation algebraically using the Fundamental Theorem of Algebra, the Rational Root Theorem, Descartes' Rule of Signs, and the Factor Theorem. (Hint: If the numbers are large, graph the function first using GeoGebra to help you find one of the zeros. Use that zero to find the depressed equation which can be solved by factoring or the quadratic formula.) Substitute 0 for the function notation and, using graphing technology, graph the function.

OpenStudy (anonymous):

Alright. Where are you confused here?

OpenStudy (anonymous):

You know the volume right? 160 right? The formula for volume is lwh. It says you however supposedly DON'T know the length, but do know the relations. The height is 2 inches bigger than the length. The width is 6 inches smaller than the length. 160 = x(x - 6)(x + 2)

OpenStudy (anonymous):

So now what?

OpenStudy (anonymous):

Multiply x(x - 6)(x + 2) out.

OpenStudy (anonymous):

So (x^2-6x)(x^2+2x) or is there more to it?

OpenStudy (anonymous):

160 = x(x - 6)(x + 2) 160 = x(x^2 - 4x - 12) 160 = x^3 - 4x^2 - 12x 0 = x^3 - 4x^2 - 12x - 160 Can you solve now using synthetic division, rational root theorem, and Descartes Rule of Signs.

OpenStudy (anonymous):

I know how to do synthetic division but what number do i use in the little box?

OpenStudy (anonymous):